# Review on the decoding algorithms for surface codes

### v4

@@ 5.1.1 Complexity
As described before, the most critical part of the MWPM decoder consists in finding the perfect matching with minimum weight once the subgraphs are constructed from the syndrome information. An algorithm to efficiently solve such computational problem was proposed by Jack Edmonds back in the 60s, the so called Blossom’s algorithm [117]. In general, the MWPM decoder is dominated by the Blossom step of the algorithm whose worst-case complexity in the number of nodes NN is $O(N^{3}\log{(N)})$ [57, 120]. However, the expected runtime of the decoder is roughly $O(N^{2})$ whenever the decoder is implemented such that all Dijktra searches are needed for computing the subgraphs where the matching of interest is needed [120, 57]. Therefore, and due to the importance speed for real-time decoding, several implementations of the MWPM decoder have been proposed such as Fowler’s implementation with $O(1)$ parallel expected runtime [127] or the more recent Sparse Blossom by Higgot and Gidney with an observed complexity of $O(N^{1.32})$ [57] and Fusion Blossom by Wu with $O(N)$, i.e. linear complexity [118, 119, 130]. Each of the implementations have their own advantages and disadvantages, as for example, Sparse Blossom has a faster single-thread performance than Fusion Blossom, but the latter supports multi-thread execution, implying that it can be faster than the former if enough cores are available [57]. Proposing faster MWPM implementations is an arduous but significant task, since large distances are precised in order to have a fault-tolerant quantum computer, and the decoding schemes also need to be scalable in the sense that they are fast enough when the distance of the code increases.
Following the previous discussions, it can be seen that the MWPM decoder follows the QMLD decoding rule as it aims to estimate the most probable error for the given syndrome. Note that, here, finding a perfect matching with minimum weight in the subgraphs formed with the measured syndrome implies that the Pauli element estimated will be the most probable to occur considering pure $X$ and $Z$ noise  [^17] This occurs because usually a i.i.d. model is assumed, implying that higher weigths imply less probability.. Applying the suboptimal decoding rule has been observed to be an efficient for low physical error probabilities [120].
@@ 5.1.2 Performance and threshold
In FIG. 8 we plot the performance of the rotated planar code in terms of the logical error probability ($P_{L}$), that is, the probability of the decoding process failing in predicting an error given its syndrome, as a function of the physical error probability ($p=p_{x}+p_{y}+p_{z}$) whenever it is decoded using a MWPM decoder. Two noise models are considered: the top figure considers an i.i.d. depolarizing error model ($p_{x}=p_{y}=p_{z}$), while in the bottom figure considers an i.i.d. biased Pauli channel with bias $\eta=100$. The results show how the MWPM decoder performs better when considering noise channels closer to the depolarizing channel. Specifically, not only the logical error probabilities are significantly higher for the biased case when a physical error probability is fixed, but also the probability threshold $p_{th}$ is lower. This performance decrease when the channel is biased can be explained by the fact that both subgraphs are considered independently. Considering a bias towards $Z$-noise results in the $Z$-subgraph correspondent to the $Z$-checks being more dense, i.e. more non-trivial syndromes are triggered, as opposed to the $X$-checks one. This results in the $Z$-subgraph reaching the probability threshold before the total physical error probability reaches the threshold of the depolarizing channel. Further increasing the bias of the channel will produce a decrease of the $p_{th}$ until the extreme value of $\eta\rightarrow\infty$, that is, a pure dephasing error model. At such point, all triggered syndromes will correspond to the same subgraph, i.e., the right column in FIG. 6. Table 1, shows some $p_{th}$ for different biases when decoded using the standard MWPM decoder for the rotated planar code.
Figure 7: Two $Z$-error chains that share the same endpoints and the stabilizing operator checks which separate them.
@@ 5.1.3 Modifications and re-weighting for specific noise models
Biased channels are important since experimentally implemented qubits have shown lower dephasing times, $T_{2}$, than relaxation times, $T_{1}$, which implies that those qubits are more prone to suffer from dephasing errors than bit-flip errors (recall the GAPDPTA channel) [30, 35, 104]. Bias values in the range of $\eta\in[1,10^{6}]$ are typical depending on the technology used for constructing the qubits [30, 78]. Therefore, the way to deal with biased noise has been studied by modifying the surface code structure so that the $Z$-subgraph contains more information than the $X$-subgrahp, i.e. rectangular surface codes, [41, 87, 131] or by modifying the MWPM decoder by making it aware of the symmetries of the code [132]. In addition, the noise in experimentally constructed hardware is not identically distributed (recall the i.ni.d. error model) [88, 104, 133, 106]. The performance of surface codes is significantly affected by such non-uniformity of the noise in the data qubits of the lattice, as some of the data qubits will have a higher tendency to suffer from errors than other, and the standard MWPM decoder calculates the perfect matching considering that all the qubits are equiprobable. Thus, methods based on reweighting the syndrome subgraphs as a function of the probability of error of each of the qubits have been proposed so that the MWPM problem is solved over a weighted graph that takes such effects in consideration [88, 104]. Another limitation of the standard MWPM decoder is the fact that since the $Z$ and $X$-subgraphs are decoded independently, $Y$-errors are underestimated. In fact, since $Y$-errors are combinations of bit- and phase-flips, it can be considered that there exists a correlation among them. Therefore, information can be passed from the $Z$-subgraph to the $X$-subgraph, and vice versa, so that the correlation betwen those events is taken into account by reweighting the other subgraph as a function of what has been estimated in the other [88, 134, 135].
@@ 5.1.4 Measurement errors
Another important thing to consider when decoding surface codes is the fact that the measured syndromes might be erroneous [32, 87, 107]. This means that the measured syndrome does not correspond to the Pauli operator affecting the data qubits after measurement. This may come due to the fact that the measurement operations are not perfect (recall SPAM errors) or due to the so called error propagation. Error propagation refers to the fact that Pauli errors on one qubit can propagate to other qubit when performing a two-qubit gate. For example, if we have an $X$-operator in the control qubit of a $\mathrm{CNOT}$ gate, such operator propagates to the target qubit, i.e. $\mathrm{CNOT}(\mathrm{X}\otimes\mathrm{I})=\mathrm{X}\otimes\mathrm{X}$. As a consequence of this, the circuit-level noise coming from the measurement qubits can propagate to the data qubits, changing the Pauli error due to propagation and, thus, making the measured error to be imperfect even with perfect SPAM. Those erroneous measurements have a big impact in code performance, lowering the code threshold in significant manner when circuit-level noise is considered [32, 87, 107, 136]. In order to deal with this problem, several measurement rounds are done before decoding so that a space-time like graph is obtained in which the MWPM problem is performed to estimate the error. Usually, $d$ measurements are recorded for a distance-$d$ surface code [32, 136]. It is noteworthy to say that once the measurements are done, a non-trivial syndrome element for a measurement that follows another one will be a measurement that has flipped from such last round, refer to the Appendix B for a more detailed description. Then, the complete graph is constructed by connecting all those non-trivial elements both spatially and temporally. This consideration enlarges the size of the graph where the perfect matching with minimum weight must be computed and, therefore, the complexity of the algorithm increases considerably. By considering this space-time decoding, the performance of the code will improve when compared to single-round decoding, but the code threshold significantly decreases compared to the perfect measurement scenario [32, 87, 107].
@@ 5.2 Union-Find decoder
The Union-Find decoder (UF) is a decoding scheme proposed by Nicolas Delfosse and Naomi Nickerson in 2017 which also consists in mapping the syndrome into a graph problem [109]. However, this decoder is based on clustering the non-trivial syndrome elements of the subgraphs by considering that Pauli errors at a known location can be treated as erasure errors. As mentioned in the error model section, an erasure error within a physical qubit can be treated as the qubit itself being in a mixed state (subjected to a random Pauli error). Therefore, having a uniform probability distribution for $I$-, $X$-, $Y$- or $Z$-operators [53] and, most importantly, a known location. Thus, for a pure erasure error model, all the qubits undergoing said erasure errors are localized. The surface code under the erasure channel can be efficiently decoded in linear time, through a method named peeling decoding scheme [137]. In light of this, the UF decoder is based on the idea of transforming the decoding problem of a surface code experiencing Pauli noise into an erasure error decoding problem. By doing this, the UF decoder achieves a decoding complexity of almost linear time $O(n\alpha(n))$ [109], where $\alpha$ is the inverse of the Ackerman function and for all practical purposes $\alpha(n)\leq 3$ [138]. In order to do so, the UF decoding process consists of two different steps: syndrome validation and erasure decoder. The syndrome validation step consists on mapping the set of Pauli errors or a mixture of Pauli and erasure errors into clusters of overall erasure errors [109]. Note that mixtures of Pauli and erasure errors can be considered by the UF decoder since the idea is to only have erasure errors. Once the step is completed, the erasure decoder is based on the peeling decoder [137].
@@ 5.2.1 Syndrome validation
FIG. 9 presents an example of syndrome validation for the $X$-check graph (note that the execution of the Z-check graph will be performed in the same manner). The top figure represents the error considered and the triggered checks. Note that we omit the $Z$-checks which we represent with orange circles for simplicity. On the second picture from the top, the clusters increase reaching the adjacent checks from the adjacent data qubits from the initial non-trivial checks. The leftmost and rightmost non-trivial checks reach the boundary and, thus, freeze, as shown in the third row. We depict frozen clusters with the cyan color. Moreover, the two triggered checks on the bottom right of the surface code also freeze as a result of the even parity of the cluster. Therefore, by the third figure, only one cluster will continue to grow. For that reason, as it can be seen in the fourth figure, the cluster grows again making contact with one of the leftmost frozen clusters. Consequently, they merge into a single cluster which is frozen because the new cluster has reached a virtual qubit. This makes the syndrome validation step to conclude, as it can be seen in the last row of FIG. 9.
Figure 9: Graphical representation of the cluster growth in a distance-9 rotated planar code under a specific error. Violet lines indicate growing clusters and cyan lines indicate parity even and frozen clusters.
@@ 5.2.2 Erasure decoder
Once the syndrome validation has been computed, the Pauli error within a surface code can be treated as an erasure error (due to its known location) and, thus, can be decoded through the peeling decoder [137]. First of all, the structure of the clusters must be that of a spanning tree in order to execute such method. In graph theory, a tree is an undirected graph in which two vertices are connected by exactly one path, and so there are no cycles. A spanning tree is a tree which contains all vertices within a graph [117]. Consequently, since the clusters after syndrome validation may have cycles, one of the associated spanning trees must be chosen. If a cluster spans from one open boundary, i.e. a boundary with virtual checks, of the surface code to the other one, this is also considered a cycle, and so it must be split in two spanning trees, one adjacent to each side of the open boundary. The vertices of degree 1 within the spanning tree, that is, the ones that are adjacent to a single edge, are named leaves [117]. For the peeling decoder, one of the leaves of each spanning tree is selected as the root of the tree. If a spanning tree contains a number of virtual qubit leaves, one of them will be considered the root. The decoding process commences by selecting a non-root leaf for each cluster and applying the following rule [109]:
@@
In FIG. 10, a possible peeling process for the error after syndrome validation in FIG. 9 is presented. Given the even parity clusters from FIG. 9, a set of four spanning trees are chosen (top left figure). In the top right figure, these spanning trees are shown with identifying colors, green edges are edges incident to leaves and brown edges represent the trunks of the trees. Moreover, the arrows indicate the growth direction from the tree root to the leaves, the peeling should be done in the opposed direction. In the second-top left image, the result of peeling all the leaves from the previous image is shown. Since most of the leaves were trivial checks, no matchings arise except for one in the bottom right spanning tree, which is denoted with a blue line. The remaining figures show the progress of the peeling decoder until reaching the bottom figure which shows the recovered error. Edges which are part of the recovered error by the peeling decoder are represented with blue lines. Notice that the recovered Pauli error in FIG. 10 is the same one as the one in FIG. 9 up to a stabilizer from the top-right $X$-Pauli error implying a successful decoding round.
Figure 10: Graphical representation of the forest peeling from a spanning forest chosen from the set of erasure errors from FIG.9.
@@ 5.3 Belief Propagation
Belief Propagation (BP) is a message-passing algorithm that can be used to solve inference problems on probabilistic graphical models [141]. It is also sometimes referred to as the Sum-Product Algorithm (SPA) [142, 143], a more general-purpose algorithm that computes marginal functions associated with a global function. Although the terms BP and SPA are essentially interchangeable, throughout this paper we will use BP to refer to the algorithm employed to decode error correction codes.
Error correction codes, irrespective of being applied in a quantum or classical paradigm, can be represented by bipartite graphs known as factor graphs. A factor graph $G=(N,E)$ is defined by a set of nodes $N=V\cup C$, where $V,C$ represent two distinct types of nodes known as variable and check nodes [^18] In the context of classical codes, variable nodes represent bits (the columns of the Parity Check Matrix (PCM)) and check nodes represent parity check operations (the rows of the PCM). The same holds true for quantum codes, but instead of representing bits and parity checks, variable nodes represent qubits and check nodes represent the action of stabilizer generators. In both paradigms, edges between variable and check nodes exist if the associated entry in the PCM is non-zero., respectively, and a set of edges $E$. Against this backdrop, BP can be used to approximate the problem of Maximum Likelihood Decoding (MLD) by exchanging messages over the factor graph representation of an error correction code. If BP runs over a graph that is a tree, it will converge to the exact MLD solution in a time bounded by the tree’s depth. In scenarios where this does not hold, i.e., when the algorithm is run over a loopy factor graph and convergence cannot be achieved, BP has proven to be a good heuristic decoding method, especially when the typical size of the loops in the graph is large.

### v5

@@ 5.1.1 Complexity
As described before, the most critical part of the MWPM decoder consists in finding the perfect matching with minimum weight once the subgraphs are constructed from the syndrome information. An algorithm to efficiently solve such computational problem was proposed by Jack Edmonds back in the 60s, the so called blossom algorithm [95]. In general, the MWPM decoder is dominated by the blossom step of the algorithm whose worst-case complexity in the number of nodes NN is $\mathcal{O}(N^{3}\log{(N)})$ [44, 97]. However, the expected runtime of the decoder is roughly $\mathcal{O}(N^{2})$ whenever the decoder is implemented such that all Dijktra searches are needed for computing the subgraphs where the matching of interest is needed [44, 97]. Therefore, and due to the importance speed for real-time decoding, several implementations of the MWPM decoder have been proposed such as Fowler’s implementation with $\mathcal{O}(1)$ parallel expected runtime [103] or the more recent sparse blossom by Higgot and Gidney with an observed complexity of $\mathcal{O}(N^{1.32})$ [44] and Fusion Blossom by Wu with $\mathcal{O}(N)$, i.e. linear complexity [96, 104]. Each of the implementations have their own advantages and disadvantages, as for example, sparse blossom has a faster single-thread performance than Fusion Blossom, but the latter supports multi-thread execution, implying that it can be faster than the former if enough cores are available [44]. Proposing faster MWPM implementations is an arduous but significant task, since large distances are precised in order to have a fault-tolerant quantum computer, and the decoding schemes also need to be scalable in the sense that they are fast enough when the distance of the code increases.
Following the previous discussions, it can be seen that the MWPM decoder follows the QMLD decoding rule as it aims to estimate the most probable error for the given syndrome. Note that, here, finding a perfect matching with minimum weight in the subgraphs formed with the measured syndrome implies that the Pauli element estimated will be the most probable to occur considering pure $X$ and $Z$ noise  [^23] This occurs because usually a i.i.d. model is assumed, implying that higher weigths imply less probability.. Applying the suboptimal decoding rule has been observed to be an efficient approach for low physical error probabilities [97].
@@ 5.1.2 Performance and threshold
In FIG. 8 we plot the performance of the rotated planar code in terms of the logical error probability ($P_{L}$), that is, the probability of the decoding process failing in predicting an error given its syndrome, as a function of the physical error probability ($p=p_{x}+p_{y}+p_{z}$) whenever it is decoded using a MWPM decoder. Two noise models are considered: the top figure considers an i.i.d. depolarizing error model ($p_{x}=p_{y}=p_{z}$), while in the bottom figure considers an i.i.d. biased Pauli channel with bias $\eta=100$. The results show how the MWPM decoder performs better when considering noise channels closer to the depolarizing channel. Specifically, not only the logical error probabilities are significantly higher for the biased case when a physical error probability is fixed, but also the probability threshold $p_{th}$ is lower. This performance decrease when the channel is biased can be explained by the fact that both subgraphs are considered independently. Considering a bias towards $Z$-noise results in the $Z$ decoding subgraph being more dense, i.e. more non-trivial syndromes are triggered, as opposed to the $X$-checks one. Having more non-trivial checks within the $Z$ decoding subgraph makes it more probable for the recovered minimum weight perfect matching not to capture the error that the code underwent, as the performance of the MWPM decreases as the physical error rate $p$ increases. This results in the $Z$-subgraph reaching the probability threshold before the total physical error probability reaches the threshold of the depolarizing channel. Further increasing the bias of the channel will produce a decrease of the $p_{th}$ until the extreme value of $\eta\rightarrow\infty$, that is, a pure dephasing error model. At such point, all triggered syndromes will correspond to the same subgraph, i.e., the right column in FIG. 6. Table 1, shows some $p_{th}$ for different biases when decoded using the standard MWPM decoder for the rotated planar code.
Figure 7: Two $Z$-error chains that share the same endpoints and the stabilizing operator checks which separate them. Since both chains are equal up to an stabilizer, they represent the same logical operation.
@@ 5.1.3 Modifications and re-weighting for specific noise models
Biased channels are important since for some experimentally implemented qubits, such as ion traps or spin qubits, have shown lower dephasing times, $T_{2}$, than relaxation times, $T_{1}$, which implies that those qubits are more prone to suffer from dephasing errors than bit-flip errors (recall Section 4) [27, 29, 83]. Bias values in the range of $\eta\in[1,10^{6}]$ are typical depending on the technology used for constructing the qubits [27, 58]. Therefore, some studies have been directed towards better adapting surface codes to $Z$ biased noise, either by introducing more $Z$-checks producing rectangular surface codes [32, 65, 106], tailoring the codes so as to have all checks susceptible to $Z$ noise [32, 65] or by modifying the MWPM decoder by making it aware of the symmetries of the code [107]. In addition, the noise in experimentally constructed hardware is not identically distributed (recall the i.ni.d. error model) [66, 83, 86, 108]. The performance of surface codes is significantly affected by such non-uniformity of the noise in the data qubits of the lattice, as some of the data qubits will have a higher tendency to suffer from errors than other, and the standard MWPM decoder calculates the perfect matching considering that all the qubits are equiprobable. Thus, methods based on reweighting the syndrome subgraphs as a function of the probability of error of each of the qubits have been proposed so that the MWPM problem is solved over a weighted graph that takes such effects in consideration [66, 83]. Another limitation of the standard MWPM decoder is the fact that since the $Z$ and $X$-subgraphs are decoded independently, $Y$-errors are underestimated. In fact, since $Y$-errors are combinations of bit- and phase-flips, it can be considered that there exists a correlation among them. Therefore, information can be passed from the $Z$-subgraph to the $X$-subgraph, and vice versa, so that the correlation between those events is taken into account by reweighting the other subgraph as a function of what has been estimated in the other [66, 109, 110].
@@ 5.1.4 Measurement errors
Another important thing to consider when decoding surface codes is the fact that the measured syndromes might be erroneous [25, 65, 87]. This means that the measured syndrome does not correspond to the Pauli operator affecting the data qubits after measurement. This may come due to the fact that the measurement operations are not perfect (recall SPAM errors) or due to the so called error propagation. Error propagation refers to the fact that Pauli errors on one qubit can propagate to other qubit when performing a two-qubit gate. For example, if we have an $X$ operator in the control qubit of a $\mathrm{CNOT}$ gate, such operator propagates to the target qubit, i.e. $\mathrm{CNOT}(\mathrm{X}\otimes\mathrm{I})=\mathrm{X}\otimes\mathrm{X}$. As a consequence of this, the circuit-level noise coming from the measurement qubits can propagate to the data qubits, changing the Pauli error due to propagation and, thus, making the measured error to be imperfect even with perfect SPAM. Those erroneous measurements have a big impact in code performance, lowering the code threshold in significant manner when circuit-level noise is considered [25, 65, 87, 111]. In order to deal with this problem, several measurement rounds are done before decoding so that a space-time like graph is obtained in which the MWPM problem is performed to estimate the error. Usually, $d$ measurements are recorded for a distance-$d$ surface code [25, 111]. It is noteworthy to say that once the measurements are done, a non-trivial syndrome element for a measurement that follows another one will be a measurement that has flipped from such last round, refer to the Appendix B for a more detailed description. Then, the complete syndrome graph is constructed by connecting all those non-trivial elements both spatially and temporally. This consideration enlarges the size of the syndrome graph where the perfect matching with minimum weight must be computed and, therefore, the complexity of the algorithm increases considerably. Measurement errors highly compromise the decoding process under single-round syndrome extraction, due to the fact that they are not accounted for and will most likely cause a logical error. For example, consider an error consisting of a single measurement error in a check, which becomes non-trivial. As has been seen in this section, the MWPM decoder will consider this check altogether with its closest virtual check and return the minimum weight perfect matching of the decoding graph as the recovered error, which will be non-trivial. By considering this space-time decoding, the performance of the code will improve when compared to single-round decoding due to the fact that measurement errors will be efficiently considered. Nevertheless, the code threshold significantly decreases compared to the perfect measurement scenario [25, 65, 87].
@@ 5.2 Union-Find decoder
The Union-Find decoder (UF) is a decoding scheme proposed by Nicolas Delfosse and Naomi Nickerson in 2017 which also consists in mapping the syndrome into a graph problem [89]. However, this decoder is based on clustering the non-trivial syndrome elements of the subgraphs by considering that Pauli errors at a known location can be treated as erasure errors. As mentioned in the error model section, an erasure error within a physical qubit can be treated as the qubit itself being in a mixed state (subjected to a random Pauli error). Therefore, having a uniform probability distribution for $I$-, $X$, $Y$ or $Z$ operators and, most importantly, a known location. Thus, for a pure erasure error model, all the qubits undergoing said erasure errors are localized. The surface codes under the erasure channel can be efficiently decoded in linear time, through a method named peeling decoding scheme [112]. In light of this, the UF decoder is based on the idea of transforming the decoding problem of a surface code experiencing Pauli noise into an erasure error decoding problem. By doing this, the UF decoder achieves a decoding complexity of almost linear time $\mathcal{O}(n\alpha(n))$ [89], where $\alpha$ is the inverse of the Ackerman function and for all practical purposes $\alpha(n)\leq 3$ [113]. In order to do so, the UF decoding process consists of two different steps: syndrome validation and erasure decoder. The syndrome validation step consists on mapping the set of Pauli errors or a mixture of Pauli and erasure errors into clusters of overall erasure errors [89]. Note that mixtures of Pauli and erasure errors can be considered by the UF decoder since the idea is to only have erasure errors. Once the step is completed, the erasure decoder is based on the peeling decoder [112].
@@ 5.2.1 Syndrome validation
FIG. 9 presents an example of syndrome validation for the $X$-check graph (note that the execution of the Z-check graph will be performed in the same manner). FIG. 9 a) represents the error considered and the triggered checks. On FIG. 9 b), the clusters increase reaching the adjacent checks from the adjacent data qubits from the initial non-trivial checks. The leftmost and rightmost non-trivial checks reach the boundary and, thus, freeze, as shown in the third row. We depict frozen clusters with the cyan color. Moreover, the two triggered checks on the bottom right of the rotated code also freeze as a result of the even parity of the cluster. Therefore, by FIG. 9 c), only one cluster will continue to grow. For that reason, as it can be seen in FIG. 9 d), the cluster grows again making contact with one of the leftmost frozen clusters. Consequently, they merge into a single cluster which is frozen because the new cluster has reached a virtual qubit. This makes the syndrome validation step to conclude, as it can be seen in the FIG. 9 e).
Figure 9: Graphical representation of the cluster growth in a distance-9 rotated planar code under a specific error. In a), an error and its syndrome is presented. From b) to e) illustrate the syndrome validation process for that syndrome. Violet lines indicate growing clusters and cyan lines indicate parity even and frozen clusters.
@@ 5.2.2 Erasure decoder
Once the syndrome validation has been computed, the Pauli error within a surface code can be treated as an erasure error (due to its known location) and, thus, can be decoded with an erasure decoder. Due to its good performance and linear complexity, the erasure decoder of choice for this step is the peeling decoder [112], which will be covered in continuation. First of all, the structure of the clusters must be that of a spanning tree in order to execute such method. In graph theory, a tree is an undirected graph in which two vertices are connected by exactly one path, and so there are no cycles. A spanning tree is a tree which contains all vertices within a graph [95]. Consequently, since the clusters after syndrome validation may have cycles, one of the associated spanning trees must be chosen. If a cluster spans from one open boundary, i.e. a boundary with virtual checks, of the surface code to the other one, this is also considered a cycle, and so it must be split in two spanning trees, one adjacent to each side of the open boundary. The vertices of degree 1 within the spanning tree, that is, the ones that are adjacent to a single edge, are named leaves [95]. For the peeling decoder, one of the leaves of each spanning tree is selected as the root of the tree. If a spanning tree contains a number of virtual qubit leaves, one of them will be considered the root. The decoding process commences by selecting a non-root leaf for each cluster and applying the following rule [89]:
@@
In FIG. 10, a possible peeling process for the error after syndrome validation in FIG. 9 is presented. Given the even parity clusters from FIG. 9, a set of four spanning trees are chosen as portrayed in FIG. 10 a). In FIG. 10 b), these spanning trees are shown with identifying colors, green edges are edges incident to leaves and brown edges represent the trunks of the trees. Moreover, the arrows indicate the growth direction from the tree root to the leaves, the peeling should be done in the opposed direction. In FIG. 10 c), the result of peeling all the leaves from the FIG. 10 b) is shown. Since most of the leaves were trivial checks, no matchings arise except for one in the bottom right spanning tree, which is denoted with a blue line. The remaining FIG. 10 d), e), f), g) and h) show the progress of the peeling decoder until reaching the bottom figure which shows the recovered error. Edges which are part of the recovered error by the peeling decoder are represented with blue lines. Notice that the recovered Pauli error in FIG. 10 i) is the same one as the one in FIG. 9 a) up to a stabilizer from the top-right $X$-Pauli error implying a successful decoding round.
Figure 10: Graphical representation of the forest peeling from a spanning forest chosen from the set of erasure errors from FIG. 9. a) presents the chosen spanning forest and b) colors the spanning tree edges in green if they are leaves and brown otherwise. The peeling process lasts from c) to h), blue edges denote the non-trivial edges for the recovered error. In i), the recovered error is presented on the initial syndrome.
@@ 5.3 Belief Propagation
Belief Propagation (BP) is a message-passing algorithm that can be used to solve inference problems on probabilistic graphical models [116]. It is also sometimes referred to as the Sum-Product Algorithm (SPA) [117], which is a more general-purpose algorithm that computes marginal functions associated with a global function. Although the terms BP and SPA are essentially interchangeable, throughout this paper we will use BP to refer to the algorithm applied to decode error correction codes.
Error correction codes, irrespective of being applied in a quantum or classical paradigm, can be represented by bipartite graphs known as factor graphs. A factor graph $G=(N_{G},E_{G})$ is defined by a set of nodes $N_{G}=V\cup C$, where $V,C$ represent two distinct types of nodes known as variable and check nodes, respectively, and $E_{G}$ is a set of edges. In the context of error correction codes, factor graphs are generally referred to as Tanner graphs. Throughout the rest of this paper, we will use the terms factor graph and Tanner graph interchangeably to refer to the graphical representation of an error correction code. A Tanner graph represents the structure of the error correcting code by means of variable nodes representing the physical bits and check nodes representing the checks. Every check will be connected with edges to the physical bits that can change its parity. Moreover, the Tanner graph can be represented through a binary matrix named the parity check matrix. Within the parity check matrix, every column represents a physical bit and every row represents a check. Non-trivial elements within the parity check matrix define edges connecting the physical bit from the column to the check from the row. Then, given a binary error vector $\bar{e}$, where non-trivial elements indicate bit-flips, one can analytically obtain the syndrome as $\bar{s}=H\cdot\bar{e}\mod 2$, where $\bar{s}$ is the syndrome and $H$ the parity check matrix.
One can extend this logic to quantum error correcting codes, although taking into account some caveats [^24] For a more detailed description to as to why parity checks are described like this in the quantum realm refer to [54].. In general terms, quantum parity check matrices have $2n$ columns, due to the fact that each qubit is considered to have two parities, the $X$-error parity, which corresponds to the first $n$ columns, and the $Z$-error parity, which corresponds to the second $n$ columns. A $Y$-error is the result of a non-trivial $X$ and $Z$-error parity within the same qubit. Moreover, the checks are also represented through the rows. As an example, we can study the distance-$3$ planar code, which is defined by the set of stabilizer generators shown in (9) (relisted below to aid the reader):
$s_{1}$ $=Z_{2}Z_{3},$
$s_{2}$ $=Z_{1}Z_{2}Z_{4}Z_{5},$
$s_{3}$ $=Z_{5}Z_{6}Z_{8}Z_{9},$
$s_{4}$ $=Z_{7}Z_{8},$
$s_{5}$ $=X_{1}X_{4},$
$s_{6}$ $=X_{4}X_{5}X_{7}X_{8},$
$s_{7}$ $=X_{2}X_{3}X_{5}X_{6},$
$s_{8}$ $=X_{6}X_{9},$
As expressed in previous sections, the rotated code is CSS, which allows us to consider two separate parity check matrices, one for the checks susceptible to $X$-errors and another for the checks susceptible to $Z$-errors. Therefore, the parity check matrix susceptible to $X$-errors is given by stabilizers $s_{1},s_{2},s_{3}$ and $s_{4}$ and has the following expression:
$H_{X}=\begin{pmatrix}0&1&1&0&0&0&0&0&0\\ 1&1&0&1&1&0&0&0&0\\ 0&0&0&0&1&1&0&1&1\\ 0&0&0&0&0&0&1&1&0\\ \end{pmatrix}$, (21)
and the expression for the parity check matrix susceptible to $Z$-errors is given by stabilizers $s_{1},s_{2},s_{3}$ and $s_{4}$ and takes the form:
$H_{Z}=\begin{pmatrix}1&0&0&1&0&0&0&0&0\\ 0&0&0&1&1&0&1&1&0\\ 0&1&1&0&1&1&0&0&0\\ 0&0&0&0&0&1&0&0&1\\ \end{pmatrix}$. (22)
Within both $H_{X}$ and $H_{Z}$, the columns represent the $X$ and $Z$-parity of the data qubits respectively and the rows represent the stabilizer generators.
Moreover, the CSS structure of the rotated planar code allows us to compute syndromes as: $\bar{s}_{X}=H_{X}\cdot\bar{e}_{X}\mod 2$ and $\bar{s}_{Z}=H_{Z}\cdot\bar{e}_{Z}\mod 2$, where , $\bar{e}_{X}$ ($\bar{e}_{Z}$) is a binary error vector correspondent to the bit-flips (phase-flips) and $\bar{s}_{X}$ ($\bar{s}_{Z}$) is a syndrome for the $X$ ($Z$) parity. Moreover, from $H_{X}$ and $H_{Z}$, we construct the Tanner graph as follows:
  1. 1.
    For every data qubit in the code, add two variable nodes to the graph. Place these nodes at the center of the graph. Our example will have $9$-pairs of variable nodes. Each node represents a different error parity, nodes to the left indicate $X$-parity and nodes to the right indicate $Z$-parity.
  2. 2.
    For every stabilizer generator in the code, add a check node to the graph. For checks correspondent to $H_{X}$, place the check nodes to the left of the variable nodes, checks correspondent to $H_{Z}$ should be placed to the right of the variable nodes.
  3. 3.
    Add an edge from every stabilizer generator to the parity it is susceptible to. This corresponds to the non-zero elements for their row in their parity check matrix. For example, $s_{1}$ corresponds to the first row in $H_{X}$, therefore it should be connected to the variable nodes $2$ and $3$.
In FIG. 12, the distance-$3$ rotated planar code with parity check matrices (21) and (22) is illustrated altogether with its Tanner graph representation. Consider that, as has been seen, every data qubit is portrayed through two parities, the $X$-parity and the $Z$-parity.This is further represented in the Tanner graph of FIG. 12, where we can see each qubit represented by two variable nodes, one interacting with $X$-checks and another interacting with $Z$-checks.
Figure 12: On the top, a representation of a distance-3 rotated planar code where both data and check qubits are labelled. On the bottom, the Tanner graph of said code. Note that the Tanner graph represents which data qubits interact with which checks.
Against this backdrop, BP can be understood as an algorithm that approximates the marginal probabilities of undergoing a parity change for every variable node in the Tanner graph through message exchanging between nodes. If BP runs over a graph that is a tree, it will converge to the exact marginal probability solution in a time bounded by the tree’s depth. That is, the marginal probabilities will provide an error which will be consistent to the syndrome. In scenarios where this does not hold, as is the case with most factor graphs that represent error correction codes, BP has proven to be a good heuristic decoding method, especially when the typical size of the loops in the graph is large.